# A Novel Simulation Approach for Evaporation Processes in Small Gaps of Industrial Applications

### v1

@@ 1 Introduction
In 1805 Young formulated a model to describe the change in pressure $\Delta p$ at a fluid-fluid interface, known as the capillary pressure, which depends on the interface curvature $\kappa$ [5]. In the same year, this model was mathematically described by Laplace yielding the Young-Laplace equation [6]. Concus and Finn studied the shape of liquid surfaces in a wedge and showed that the Young-Laplace equation only has a bounded solution if $\theta>\pi/2-\alpha$, where $\theta$ is the contact angle as depicted in Figure 1 and $\alpha$ is the half wedge angle [7].
@@
Two centuries after Stefan’s findings, Camassel et al. proposed a simplified analytical model for square capillaries that neglects gravity and viscous effects to calculate the evaporation rate [3]. A few years later, Yiotis et al. developed an analytical model that also takes gravity and viscous effects into account [12] to determine the three-phase drying process in non-circular capillaries which was previously proposed by Chauvet et al. [11, 9]. Since analytical solutions are mostly limited to simple geometries, such as round capillaries, numerical models were developed to enable prediction of evaporation in arbitrary shaped geometries.
A major problem in common numerical methods that enable a full resolution of the free liquid surface, e.g. Volume-of-Fluid (VoF) or Front-Tracking, is the numerical handling of the mass source term at the liquid-gas interface due to evaporation in combination with the interface movement [13, 1, 14]. Irfan and Muradoglu developed a model to incorporate concentration-driven evaporation for the front-tracking method [15]. Their model is tested for 2D cases but they state that it could be easily adapted to 3D cases as well. Nonetheless, the Front-Tracking method is difficult to parallelize, thus it’s hardly applicable to complex simulations with several Million grid cells [16]. For the VoF approach a volume fraction field is used which allows interface capturing. The evaporation mass source term results in a non-divergence-free volume fraction field which tends to produce artificial spurious currents. To counteract this unintended behavior due to jump conditions at the interface, several methods were developed to dampen spurious currents [14, 17, 18, 13]. In all previously mentioned works the Navier-Stokes equations are solved in combination with an energy equation, which results in a high computational effort. However, Schweigler et al. used an Allen-Cahn phase-field approach where only an energy functional is solved to predict evaporation behavior in capillaries [4]. Although this method yields good results for arbitrary capillary shapes, it strongly relies on numerical parameters, gathered from calibration with experimental data.
@@ 2 Methodology
where $\rho_{\mathrm{g}}$ is the gas density, $D$ is the diffusion coefficient of the evaporating species in the inert species, $Y_{\mathrm{v,sat}}$ is the saturation vapor mass fraction, $\,\text{d}Y/\text{d}\bm{n}$ is the partial derivative of vapor mass fraction with respect to the interface normal vector $\bm{n}$ and $A_{\Sigma}$ is the liquid-gas interface surface area which depends on the capillary geometry $\Omega$ and the mean diffusion length $\bar{\chi}$. The times $t_{i}$, at which the evaporation rate is evaluated, are unknown prior to the simulation. Thus, the function $f\,:\,\bar{\chi}\xrightarrow{}\dot{m}_{\mathrm{ev}}$ is solved which can be transformed into $g\,:\,m_{\mathrm{v}}\xrightarrow{}\dot{m}_{\mathrm{ev}}$ by substituting $\bar{\chi}=m_{\mathrm{v}}/(A_{\mathrm{c}}\rho_{\mathrm{l}})$ where $m_{\mathrm{v}}$ is the evaporated vapor mass. Since $\dot{m}_{\mathrm{ev}}=\,\text{d}m_{\mathrm{v}}/\text{d}t$, the time evaluates to
@@
Utilizing Equation 2, the water mass $m_{\mathrm{w}}(t_{\mathrm{ev}})$ at a certain point in time is implicitly determined as
$m_{\mathrm{w}}(t_{i})=m_{\mathrm{w}}(t_{i}=0)-m_{\mathrm{v}}(t_{i}).$ (3)
@@
    Diffusion simulation: The evaporation rate is determined using OF which solves the Laplace equation $\nabla\cdot(D\nabla Y_{v})$.
@@ 2.1 Surface formation simulation
In this work, capillaries with various cross section shapes are considered. For all cases, a parameterized input file for SE is created with the mean diffusion length as a parameter. All geometries are centered around the origin and the bottom of the capillary is aligned with the plane at $z=0$ as visualized in Figure 1.
@@
In the surface formation simulations, the liquid surface is initialized as a plane at $z=h_{\mathrm{l}}$ where $h_{\mathrm{l}}$ is the liquid filling height (see Figure 1) and then discretized as facets, edges and vertices. In order to comply with the contact angle boundary condition, the surface must be deformed. The surface deformation is done by calculating a vertex displacement vector depending on the local surface energy contribution, thus energy constraints must be implemented at the walls. The total liquid-gas surface energy at the walls is
@@
The initial filling height $h_{\mathrm{l}}$ depends on the mean diffusion length as $h_{\mathrm{l}}=h_{\mathrm{c}}-\bar{\chi}$, thus the volume at the start of simulation is given as $V_{0}=h_{\mathrm{l}}A_{\mathrm{c}}$. Since the liquid is assumed to be incompressible, its volume must be conserved. In order to comply with volume conservation, the volume error at each iteration is calculated as
@@ 2.3 Vapor diffusion simulation
interface Dirichlet $Y_{\mathrm{v,sat}}$
@@
Table 2: Numerical settings and parameters that are used for the diffusion simulations.
@@ 3 Results and discussion
Simulations are conducted for the evaporation of water in air for round and square capillaries according to the experiments from Schweigler [4]. A height of $h_{\mathrm{c}}=50$ mm is used for all capillaries as in the experiments. The temperature $T=23$C and the pressure $p=101325$ Pa are constant for all simulations. The environmental mass fraction $Y_{\mathrm{v,\infty}}$ is set according to a relative humidity of $\varphi=p_{\mathrm{v}}/p_{\mathrm{v,sat}}=17$%. The temperature dependence of density $\rho_{\mathrm{g}}$ and diffusion coefficient $D$ are incorporated using the ideal gas law for the humid air gas mixture and the Fuller method [24, 25] respectively. The thermodynamical properties of water and air under the given conditions are listed in Table 3.
@@
Environment vapor mass fraction $Y_{\mathrm{v,\infty}}$ $2.984$ $\mathrm{\,g/kg}$ Equations 11, 12 and 13
Saturated vapor mass fraction $Y_{\mathrm{v,sat}}$ $17.71$ $\mathrm{\,g/kg}$ Equations 11, 12 and 13
@@ 3.1 Round capillaries
Figure 6: Simulated fluid surface from SE for the round capillary with a diameter of $d_{\mathrm{c}}=10\mathrm{\,mm}$, height of $h_{\mathrm{c}}=50\mathrm{\,mm}$ and a contact angle of $\theta=20^{\circ}$.
@@
Figure 7: Meniscus sections of the round capillaries with $d_{\mathrm{c}}=1\mathrm{\,mm}$ (left) and $d_{\mathrm{c}}=10\mathrm{\,mm}$ (right) for various mean diffusion lengths $\bar{\chi}$ are displayed in solid colored lines. Utilizing the symmetry axis, only the right arc of the section is depicted and the sections are shifted in $z$-direction such that their deepest points collapse. The black marked lines show the meniscus approximation using a spherical cap and the Bezier method according to [27].
Figure 7 shows that the meniscus deepest point lies in the capillary center ($r/r_{\mathrm{c}}=0$) in all simulated cases. The gradient $\,\text{d}z/\,\text{d}r\xrightarrow{}0$ for $r\xrightarrow{}0$ yielding a continuously differentiable surface for $r\in(-r_{\mathrm{c}},r_{\mathrm{c}})$. Further, the water rises at the capillary walls ($r/r_{\mathrm{c}}=1$) in both cases in order to meet the contact angle boundary condition. The meniscus shape where this boundary condition is satisfied, is referred to as fully developed shape in the following. This shape corresponds to the meniscus of a motionless fluid in a round capillary of infinite height. In the R10 case, the meniscus is not fully developed for $\bar{\chi}<3\mathrm{\,mm}$, as visible in the right graph. In those simulations, the water surface reaches the top of the capillary which results in an increasing contact angle in order to comply with the volume conservation. However, in the R1 case, the meniscus is fully developed for all mean diffusion lengths due to the liquid elevation not being limited at the top of the capillary.
@@ 3.2 Square capillaries
Previous experiments by Schweigler et al. [4] show that the evaporation rate increases significantly for capillaries with a square cross-section compared to round capillaries with a similar cross-sectional area. The reason for higher evaporation rates is the presence of liquid fingers in the corners of the capillary, resulting in a local reduction of the diffusion length [4]. In order to elaborate the influence of liquid fingers on the evaporation rate, simulations of square capillaries with an inner edge length of 1mm, 4mm and $1\mathrm{\,mm},\,4\mathrm{\,mm}\text{ and }13\mathrm{\,mm}$ (cases S1, S4, S13) are performed. As a first step, the liquid surface formation is simulated in SE for the mean diffusion lengths in Equation 16. The resulting water surface for case S1 is shown as an example in Figure 9.
@@ 3.3 Temperature dependence
Since the temperature is not a direct input to the diffusion simulation, it must be used to determine the vapor mass fractions at the meniscus and environment, according to Equations 11, 12 and 13. In order to evaluate the evaporation rate, the gas mixture density $\rho_{\mathrm{g}}$ and the diffusion coefficient $D$ in Equation 15 must be adapted. The calculated values are listed in Table 4 and the respective results are plotted in Figure 12.
@@ 3.5 Generalization
To better compare the evaporation rate in various gaps, the results are normalized by the cross-sectional area of the gap. The graph clearly shows that the normalized evaporation rate is highest for $l_{\mathrm{G}}/w_{\mathrm{G}}=1$ and decreases towards a limit for higher length-to-width ratios. As described in Section 3.2, the liquid finger thickness increases only slightly for higher capillary widths which results in a decreasing influence of liquid fingers for higher $l_{\mathrm{G}}/w_{\mathrm{G}}$ ratios. The mass plot allows for a similar conclusion, as the water volume and absolute evaporation rate depend linearly on the gap cross-sectional area. With the presented method, the evaporation in arbitrary capillaries can be simulated within a reasonable simulation time, as far as the geometry can be described in SE. Future work should focus on incorporating STL (stereolithography) files in SE and conducting experiments to validate the simulated results shown here.

### v2

# A novel simulation approach for concentration-driven evaporation in capillaries

@@ Abstract
journal: International Journal of Heat and Mass Transfer
@@ 1 Introduction
In 1805, Young formulated a model to describe the change in pressure $\Delta p$ at a fluid-fluid interface, known as the capillary pressure, which depends on the interface curvature $\kappa$ [5]. In the same year, this model was mathematically described by Laplace yielding the Young-Laplace equation [6]. Concus and Finn studied the shape of liquid surfaces in a wedge and showed that the Young-Laplace equation only has a bounded solution if $\theta>\pi/2-\alpha$, where $\theta$ is the contact angle as depicted in Figure 1 and $\alpha$ is the half wedge angle [7].
$h_{\mathrm{l}}$
@@
Two centuries after Stefan’s findings, Camassel et al. proposed a simplified analytical model for square capillaries that neglects gravity and viscous effects to calculate the evaporation rate [3]. A few years later, Yiotis et al. developed an analytical model that also takes gravity and viscous effects into account [12] to determine the three-phase drying process in non-circular capillaries which was previously proposed by Chauvet et al. [11, 9]. Since analytical solutions are mostly limited to simple geometries, such as round capillaries, numerical models were developed that enable prediction of evaporation in arbitrary shaped geometries.
A major problem in common numerical methods that enable a full resolution of the free liquid surface, e.g. Volume-of-Fluid (VoF) or Front-Tracking, is the numerical handling of the mass source term at the liquid-gas interface due to evaporation in combination with the interface movement [13, 1, 14]. Irfan and Muradoglu developed a model to incorporate concentration-driven evaporation for the front-tracking method [15]. Their model is tested for 2D cases but they state that it could be easily adapted to 3D cases as well. Nonetheless, the Front-Tracking method is difficult to parallelize, thus it’s hardly applicable to complex simulations with several Million grid cells [16]. For the VoF approach a volume fraction field is used which allows interface capturing. The evaporation mass source term results in a non-divergence-free volume fraction field which tends to produce artificial spurious currents. To counteract this unintended behavior due to jump conditions at the interface, several methods were developed to dampen spurious currents [14, 17, 18, 13]. In all previously mentioned works, the Navier-Stokes equations are solved in combination with an energy equation which results in a high computational effort. However, Schweigler et al. used an Allen-Cahn phase-field approach where only an energy functional is solved to predict evaporation behavior in capillaries [4]. Although this method yields good results for arbitrary capillary shapes, it strongly relies on numerical parameters, gathered from calibration with experimental data.
@@ 2 Methodology
where $\rho_{\mathrm{g}}$ is the gas density, $D$ is the diffusion coefficient of the evaporating species in the inert species, $Y_{\mathrm{v,sat}}$ is the saturation vapor mass fraction, $\,\text{d}Y/\text{d}\bm{n}$ is the partial derivative of vapor mass fraction with respect to the interface normal vector $\bm{n}$ and $A_{\Sigma}$ is the liquid-gas interface surface area which depends on the capillary geometry $\Omega$ and the mean diffusion length $\bar{\chi}$. The latter describes the liquid filling height of a cavity in case of a flat surface (see Figure 1) and can also be considered as the distance, vapor molecules have to be transported out of the capillary. It is calculated as
$\bar{\chi}=h_{\mathrm{c}}-h_{\mathrm{l}}=\frac{m_{\mathrm{ev}}}{A_{\mathrm{c}}\rho_{\mathrm{l}}}$ (2)
with the capillary height $h_{\mathrm{c}}$ and the actual liquid filling height $h_{\mathrm{l}}$ according to Figure 1. Further, $m_{\mathrm{ev}}$ is the evaporated liquid mass, $A_{\mathrm{c}}$ is the capillary cross-section and $\rho_{\mathrm{l}}$ is the liquid density. In addition to the spatial meaning, the mean diffusion length also has a temporal meaning, since $\bar{\chi}$ increases with time as the water evaporates and the meniscus recedes inside the capillary.
The times $t_{i}$, at which the evaporation rate is evaluated, are unknown prior to the simulation. Thus, the function $f\,:\,\bar{\chi}\xrightarrow{}\dot{m}_{\mathrm{ev}}$ is solved which can be transformed into $g\,:\,m_{\mathrm{ev}}\xrightarrow{}\dot{m}_{\mathrm{ev}}$ by substituting $\bar{\chi}$ using Equation 2. Since $\dot{m}_{\mathrm{ev}}=\,\text{d}m_{\mathrm{ev}}/\text{d}t$, the time evaluates to
@@
Utilizing Equation 3, the water mass $m_{\mathrm{w}}(t_{i})$ is determined as
$m_{\mathrm{w}}(t_{i})=m_{\mathrm{w}}(t_{i}=0)-m_{\mathrm{ev}}(t_{i}).$ (4)
@@
    Diffusion simulation: The evaporation rate is determined using OF which solves the Laplace equation $\nabla\cdot(\rho_{\mathrm{g}}D\nabla Y_{v})$.
@@ 2.1 Surface formation simulation
In this work, capillaries with various cross section shapes are considered. For all cases, a parameterized input file for SE is created with the mean diffusion length $\bar{\chi}$ as an input parameter. All geometries are centered around the origin and the bottom of the capillary is aligned with the plane at $z=0$ as visualized in Figure 1.
@@
In the surface formation simulations, the liquid surface is initialized as a plane at $z=h_{\mathrm{l}}$ where $h_{\mathrm{l}}$ is the liquid filling height (see Figure 1 and C) and then discretized as facets, edges and vertices. In order to comply with the contact angle boundary condition, the surface must be deformed. The surface deformation is done by calculating a vertex displacement vector depending on the local surface energy contribution, thus energy constraints must be implemented at the walls. The total liquid-gas surface energy at the walls is
@@
The initial filling height $h_{\mathrm{l}}$ depends on the mean diffusion length as $h_{\mathrm{l}}=h_{\mathrm{c}}-\bar{\chi}$ (see Equation 2), thus the volume at the start of simulation is given as $V_{0}=h_{\mathrm{l}}A_{\mathrm{c}}$. Since the liquid is assumed to be incompressible, its volume must be conserved. In order to comply with volume conservation, the volume error at each iteration is calculated as
@@ 2.3 Vapor diffusion simulation
liquid surface Dirichlet $Y_{\mathrm{v,sat}}$
@@
Table 2: Numerical settings and parameters that are used for the diffusion simulations in OpenFOAM. These parameters are generally applicable to all cases discussed in this study.
@@ 3 Results and discussion
Simulations are conducted for the evaporation of water in air for round and square capillaries according to the experiments from Schweigler [4]. A height of $h_{\mathrm{c}}=50$ mm is used for all capillaries as in the experiments. The temperature $T=23$C and the pressure $p=101325$ Pa are constant for all simulations. The environmental mass fraction $Y_{\mathrm{v,\infty}}$ is set according to a relative humidity of $\varphi=p_{\mathrm{v}}/p_{\mathrm{v,sat}}=17$. The temperature dependence of density $\rho_{\mathrm{g}}$ and diffusion coefficient $D$ are incorporated using the ideal gas law for the humid air gas mixture and the Fuller method [24, 25] respectively. The thermodynamical properties of water and air under the given conditions are listed in Table 3.
@@
Environment vapor mass fraction $Y_{\mathrm{v,\infty}}$ $2.984$ $\mathrm{\,g/kg}$ Equations 12, 13 and 14
Saturated vapor mass fraction $Y_{\mathrm{v,sat}}$ $17.71$ $\mathrm{\,g/kg}$ Equations 12, 13 and 14
@@ 3.1 Round capillaries
Figure 6: Simulated fluid surface from SE for the round capillary with a diameter of $d_{\mathrm{c}}=1\mathrm{\,mm}$, height of $h_{\mathrm{c}}=50\mathrm{\,mm}$ and a contact angle of $\theta=20^{\circ}$.
@@
$h_{\mathrm{c}}$
$\bar{\chi}$
$h_{\mathrm{c}}$
$\bar{\chi}$
Figure 7: Meniscus sections of the round capillaries with $d_{\mathrm{c}}=1\mathrm{\,mm}$ (left) and $d_{\mathrm{c}}=10\mathrm{\,mm}$ (right) for various mean diffusion lengths $\bar{\chi}$ are displayed in solid colored lines. Utilizing symmetry, only the right arc of the section is depicted and the sections are shifted vertically to align their deepest points. Black lines show the spherical cap [27] and the Bezier approximation [28]. The upper capillary boundary constrains the meniscus shape, especially in the R10 case at smaller mean diffusion lengths which results in a different meniscus shape. This effect is less pronounced in the R1 case due to the larger mean diffusion length relative to the capillary diameter.
Figure 7 shows that the meniscus deepest point lies in the capillary center ($r/r_{\mathrm{c}}=0$) in all simulated cases. The gradient $\,\text{d}z/\,\text{d}r\xrightarrow{}0$ for $r\xrightarrow{}0$ yielding a continuously differentiable surface for $r\in(-r_{\mathrm{c}},r_{\mathrm{c}})$. Further, the water rises at the capillary walls ($r/r_{\mathrm{c}}=1$) in both cases in order to meet the contact angle boundary condition. In the R1 case, the apparent contact angle at the wall matches the specified contact angle (here $\theta=20^{\circ}$) well for all investigated values of $\bar{\chi}$. However, in the R10 case for $\bar{\chi}<3\mathrm{\,mm}$, the apparent contact angle is greater than the specified contact angle. In round capillaries of constant width and without an upper boundary, the liquid surface shape does not depend on liquid filling height or mean diffusion length respectively. However, the investigated capillaries have an upper boundary which leads to the apparent contact angle decreasing at the beginning (starting from a full capillary) until the apparent contact angle matches the specified contact angle.
@@ 3.2 Square capillaries
Previous experiments by Schweigler et al. [4] show that the evaporation rate increases significantly for capillaries with a square cross-section compared to round capillaries with a similar cross-sectional area. The reason for higher evaporation rates is the presence of liquid fingers in the corners of the capillary, resulting in a local reduction of the diffusion length [4]. In order to elaborate the influence of liquid fingers on the evaporation rate, simulations of square capillaries with an inner edge length of 1mm, 4mm and $1\mathrm{\,mm},\,4\mathrm{\,mm}\text{ and }13\mathrm{\,mm}$ (cases S1, S4, S13) are performed. As a first step, the liquid surface formation is simulated in SE for the mean diffusion lengths in Equation 17. C provides an example of the iteration process from a flat plane to a curved meniscus. The resulting water surface for the case S1 is shown as an example in Figure 9.
@@ 3.3 Temperature dependence
Since the temperature is not a direct input to the diffusion simulation, it must be used to determine the vapor mass fractions at the meniscus and environment, according to Equations 12, 13 and 14. In order to evaluate the evaporation rate, the gas mixture density $\rho_{\mathrm{g}}$ and the diffusion coefficient $D$ in Equation 16 must be adapted. The calculated values are listed in Table 4 and the respective results are plotted in Figure 12.
@@ 3.5 Generalization
To better compare the evaporation rate in various gaps, the results are normalized by the cross-sectional area $A_{\mathrm{G}}=l_{\mathrm{G}}w_{\mathrm{G}}$ of the gap. The graph clearly shows that the normalized evaporation rate is highest for $l_{\mathrm{G}}/w_{\mathrm{G}}=1$ and decreases towards a limit for higher length-to-width ratios. As described in Section 3.2, the liquid finger thickness increases only slightly for higher capillary widths which results in a decreasing influence of liquid fingers for higher $l_{\mathrm{G}}/w_{\mathrm{G}}$ ratios. The mass plot allows for a similar conclusion, as the water volume and absolute evaporation rate depend linearly on the gap cross-sectional area. With the presented method, the evaporation in arbitrary capillaries can be simulated within a reasonable simulation time, as far as the geometry can be described in SE. Future work should focus on incorporating STL (stereolithography) files in SE and conducting experiments to validate the simulated results shown here.
@@ Appendix B Reliability
Appendix C Surface Evolver Iteration Process
For the surface formation simulations the liquid surface is initialized as a plane at the height $z=h_{\mathrm{l}}=h_{\mathrm{c}}-\bar{\chi}$. The initial surface for a square capillary with a width of $d_{\mathrm{c}}=4\mathrm{\,mm}$ at $\bar{\chi}=10\mathrm{\,mm}$ or $h_{\mathrm{l}}=40\mathrm{\,mm}$ respectively, is illustrated on the left of Figure 17. Further, the figure shows the evolving surface every 50 iterations and the steady state surface on the right side.
$h_{\mathrm{c}}$
$h_{\mathrm{l}}$
$\bar{\chi}$
00
$50$
$100$
$150$
$250$
$200$
$461$
$N_{\mathrm{iter}}$
Figure 17: Liquid surface shape in a square capillary of width $d_{\mathrm{c}}=4\mathrm{\,mm}$ at different stages of iterations during the surface formation simulation in Surface Evolver. (left) Initialized surface at $\bar{\chi}=10\mathrm{\,mm}$ or $h_{\mathrm{l}}=40\mathrm{\,mm}$ respectively, (right) fully developed liquid surface after termination criteria in SE-FIT algorithm is reached.
As shown in Figure 17, the liquid fingers in the corner of the capillary evolve rapidly in the beginning until they reach the upper capillary limit at $z=h_{\mathrm{c}}$. After 250 iterations, the liquid surface only changes slightly to achieve a steady state solution. The termination criteria for a steady state solution is defined as
$\bar{E_{\gamma}}^{i}=\frac{1}{5}\sum_{j=i-4}^{i}E_{\gamma}^{j},$ (22)
$\bar{E_{\gamma}}^{i+1}-\bar{E_{\gamma}}^{i}<10^{-6}.$ (23)
where $\bar{E_{\gamma}}^{i}$ is the mean surface energy at iteration $i$ over the last 5 iterations and $E_{\gamma}^{j}$ is the surface energy in iteration $j$.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT and DeepL in order to improve the readability and perform grammar checking. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Acknowledgements
We gratefully thank Tobias Tolle for his invaluable guidance and support throughout this project. His expertise in coding and insightful supervision were instrumental in the development and implementation of the simulation methods. We especially appreciate his willingness to address our coding questions and provide helpful direction.